OpenStudy (anonymous):

If f(x)= 2x^3 -4x , find f'(3)

OpenStudy (ranga):

Find the derivative Then put x = 3 and evaluate the derivative.

OpenStudy (anonymous):

Yeah but can you please find the derivative ?

OpenStudy (ranga):

Derivative of x^n is nx^(n-1) f(x) = 2x^3 - 4x f'(x) = d/dx(2x^3 - 4x) = d/dx(2x^3) - d/dx(4x) = 2 * d/dx(x^3) - 4 * d/dx(x) = 2 * (3x^2) - 4 * (1) f'(x) = 6x^2 - 4 f'(3) = 6(3)^2 - 4 = 54 - 4 = 50